I am not quite sure where this falls into but here is the question.

rearrange this formula for the indicated variable
A=a+bh for b

Answer by lynnlo(4176)

A=a+bh
SOLVE FOR B
a=(1/2)bh
MULTIPLY BOTH SIDES BY 2
2a=bh
DIVIDE BOTH SIDES BY H
2a/h=b
B=2a/h